OpenStudy (anonymous):

What would the length of the side be if the area of a square is given (x-11)^2 ft^2

OpenStudy (campbell_st):

the area of a square is \[A = l^2\] to find the side length, take the square root of the area \[side~ length = \sqrt{l^2} ~~= l\] apply this idea to your question